OpenStudy (anonymous):

Area of a Trapezoid Solve for b1: A=1/2h(b1+b2) please show the steps thank you

jhonyy9 (jhonyy9):

A=(h(b1+b2))/2 2A=h(b1+b2) 2A/h=b1+b2 2A/h -b2=b1